OpenStudy (anonymous):

Two quadratic functions are shown. Function 1: f(x) = 2x^2 − 8x + 1 Function 2: x g(x) −2 2 −1 −3 0 2 1 17 Which function has the least minimum value and what are its coordinates? A)Function 1 has the least minimum value and its coordinates are (0, 1). B)Function 1 has the least minimum value and its coordinates are (2, −7). C)Function 2 has the least minimum value and its coordinates are (0, 2). D)Function 2 has the least minimum value and its coordinates are (−1, −3).

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

is this calculus of algebra..?

OpenStudy (anonymous):

YES IS ALGEBRA

OpenStudy (campbell_st):

ok... find the line of symmetry for function 1, the formula is \[x = \frac{-b}{2\times a}\] the function has a = 2 and b = -8 so can you just calculate that to start

OpenStudy (anonymous):

2*-8=?

OpenStudy (campbell_st):

well not quite \[x = \frac{-(-8)}{4} = 2\] so knowing the line of symmetry, you should know that this is the value of x that gives the minimum value... so substitute x = 2 into function 1 and get a value

OpenStudy (anonymous):

so, i need to solve that problem? to star?

OpenStudy (campbell_st):

no in function 1, just replace x with the number 2 and calculate the value..

OpenStudy (campbell_st):

\[y = 2 \times 2^2 - 8 \times 2 + 1\]

OpenStudy (anonymous):

-7?

OpenStudy (campbell_st):

great so the mimimum for function 1 is at the point (2, -7) now look at the table for function 2... what is the minimum y value...?

OpenStudy (anonymous):

can you give the problem and i can solve

OpenStudy (campbell_st):

no, look at the table for function 2, go down the y values and which is the minimum of smallest

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

im not good at math :/

OpenStudy (campbell_st):

the y values in the table are 2, -3, 2, 17 which is the smallest..?

OpenStudy (anonymous):

oh wait let me solve

OpenStudy (anonymous):

its 7?

OpenStudy (campbell_st):

these are the y values for function 2 y = 2, y = -3, y = 2 and y = 17 which one is the lowest..?

OpenStudy (anonymous):

-3?

OpenStudy (campbell_st):

great so in function 2, the minimum is at the point (-1, -3) y = -3 and remember function 1 has a minimum at (2, -7) y = -7 so compare the y values and decide which is lower...?

OpenStudy (anonymous):

the answer its B? IM right?

OpenStudy (campbell_st):

that's correct

OpenStudy (anonymous):

thank you so much!